OpenStudy (anonymous):

Determine the type of intersection of the two planes 3x-y+z-5=0 and 12x-4y+4z-25=0?

OpenStudy (anonymous):

Gut feeling? No intersection.

OpenStudy (anonymous):

Parallel

OpenStudy (anonymous):

no solution.

OpenStudy (anonymous):

graphs of [[z = -3*x+y]], [[z = -3*x+y+25/4]] when you solve for z and plot in 3d are parallel.